Solve for X.
(1/3+5)
-------=8
(x-1)

Answer by checkley71(8403)

(1/3+5)/(x-1)=8 or 16/3=8x-8 or 8x=16/3+8 or 8x=16/3+24/3 or 8x=40/3 or
x=40/3*1/8 or x=5/3
proof (5 1/3)/(5/3-1)=8 or (5 1/3)/(2/3)=8 or (16/3)(3/2)=8 or (16*3)/(3*2)=8 or 48/6=8 or 8=8